# One-Shot Graph Representation Learning Using Hyperdimensional Computing

### v1

@@ 3 Hyper-Dimensional Graph Learner
The proposed HD learner carries out 3 major steps:
  1. i
    Map Node features $\mathbf{x}_{v}\in\mathbb{R}^{d}$ to a intermediate HD representation $\mathbf{r}_{v}\in\{0,1\}^{\beta}$ $\forall v\in\mathcal{V}$.
  2. ii
    Construct a latent HD representation $\mathbf{z}_{v}\in\{0,1\}^{\beta}$ using $\mathbf{r}_{v}$ and $\mathcal{N}^{k}(v)$.
  3. iii
    Finally, use $\mathbf{z}_{v}$ to predict $y_{v}$
We proceed to describe each of these steps in detail.
Figure 1: An example mapping node features to a binary space using Random Hyperplanes from Efimov (2023). A point receives a value of 1 when on the left side of a hyperplane and 0 when on the right side. Utilizing multiple hyperplanes results in a binary sketch for each point.
@@ 3.1 Mapping Node Features to HD-space
The first crucial step is to map node features from $\mathbb{R}^{d}$ space to $\{0,1\}^{\beta}$ HD space such that $\beta>>d$. We do this using random hyperplane tessellations Dirksen et al. (2022). The key idea is to generate $\beta$ random hyperplanes in $\mathbb{R}^{d}$ and set the corresponding co-ordinate of the HD representation of a node based on whether the its feature vector in $\mathbb{R}^{d}$ lies on the positive or the negative side of the hyperplane. With $\beta$ such hyper-planes in hand, we can generate an $n$ dimensional binary sketch for each node. More precisely, let $\mathbf{A}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{\beta}$ be a matrix with each row $\mathbf{a}_{1}^{T},\cdots\mathbf{a}_{\beta}^{T}$ drawn from $\mathcal{N}(0,I_{d})$ and $\gamma$ be uniformly distributed on $[-\lambda,\lambda]$. For conciseness, we represent $\beta$ samples of $\gamma$ as $\Gamma\sim[-\lambda,\lambda]^{\beta}$. Using $\mathbf{A}$ and $\Gamma$, a randomized binary sketch $\mathbf{r}_{v}$ for a node $v\in\mathcal{V}$ is given by $\mathbf{r}_{v}=\text{sign}(\mathbf{A}\mathbf{x}_{v}+\Gamma)$.
@@ 3.2 Neighborhood Aggregation Scheme
Under the simplifying assumption that nodes that have identical multi-sets of features over their $k$-hop neighborhoods are likely to also have matching local topologies, we approximate the injectivity of the embedding $\mathbf{z}_{v}$ which maps each node to its HD representation by aggregating features of nodes over a $k$-hop neighborhood of $v$ as follows:
$\begin{split}\mathbf{z}_{v}=\phi\Big{(}\mathbf{r}_{v},\varphi_{1}(\{\mathbf{r}_{j};j\in\mathcal{N}^{1}(v)\}),\ldots\\ \ldots,\varphi_{K}(\{\mathbf{r}_{m};m\in\mathcal{N}^{K}(v)\})\Big{)}\end{split}$
where $\mathbf{z}_{v}$ denotes the HD embedding of node $v$ and $\phi(\cdot)$ and $\varphi_{k}(\cdot);k\in\{1,\cdots,K\}$ are injective functions.
3.2.1 Bundling for generalizing across similar nodes
We propose that for a node $v\in\mathcal{V}$; the functions $\varphi_{k}(\cdot)$ should be the Bundling operation $\oplus$ since, due to its permutation invariance, allows us to represent unordered sets (or multi-sets). Hence, we propose to use Bundling to realize $\varphi_{k}$. More precisely,
$\varphi_{k}\Big{(}\{\mathbf{r}_{j};j\in\mathcal{N}^{k}(v)\}\Big{)}=\mathop{\oplus}_{j\in\mathcal{N}^{k}(v)}\mathbf{r}_{j}$
where $\mathop{\oplus}_{j\in\mathcal{N}^{k}(v)}\mathbf{r}_{j}$ denotes binding the HD representations of all of the $k$-hop neighbors of $v$. This operation is analogous to finding a high-dimensional vector that is the most similar to all of the input HD-vectors, akin to how we find the mean in Euclidean spaces and like the mean, it is not injective. However, as noted by Xu et al. (2019), GNNs that use the mean to aggregate features over node neighborhoods Kipf and Welling (2017); Hamilton et al. (2017) perform quite well in practice because the mean operation becomes increasingly injective when graphs consist of diverse node features, which is often the case in real-world graphs.
@@ 3.2.1 Bundling for generalizing across similar nodes
3.2.2 Binding for linking node with its neighborhood
Having incorporated the Bundling operation in our framework by assigning it to the function $\varphi_{k}(\cdot)$, the remaining task involves identifying a suitable operation for the function $\phi(\cdot)$, for coupling the representation of a node obtained using bundling (see above) with the representations of nodes in its neighborhood.
@@ 3.2.2 Binding for linking node with its neighborhood
It is easy to see that if all the terms in $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ are similar, i.e., $v$ and $w$ have similar node features, and similar 1-hop and 2-hop neighborhoods, then $\mathbf{z}_{v}\approx\mathbf{z}_{w}$ as desired.
On the other hand, if the terms representing the features of the 2-hop neighbors of both $v$ and $w$ differ, it results in the dissimilarity between $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ a desired outcome. When not only the 2-hop neighbors but also the 1-hop neighbors of both $v$ and $w$ are dissimilar, the dissimilarity between $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ becomes even greater. The key property that makes binding attractive for realizing the function $\phi(\cdot)$ is its reflectivity, i.e., the distance between two HD vectors remains unchanged when both of them are bound to the same third vector.
@@
Dataset Nodes Edges Features Classes Label rate Edge density
CORA 2708 5429 1433 7 0.0563 0.0004
CiteSeer 3327 4732 3703 6 0.0569 0.0004
PubMed 19717 44338 500 3 0.0030 0.0001
Coauthor CS 18333 81894 6805 15 0.0164 0.0001
Coauthor Physics 34493 247962 8415 5 0.0029 0.0001
@@
3.2.3 Improving injectivity via Rotation
Because Binding is associative and commutative, it can adversly impact the injectivity and consequently, therepresentational power of our HD model. For example, suppose $v\in\mathcal{V}$ and $w\in\mathcal{V}$ are two nodes in a graph with encoding obtained using HD mapping that uses Bundling to realize $\varphi_{k}$ and binding to realize $\phi(\cdot)$ and $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ are computed as per Eq. (1). Consider the scenario in which $\mathbf{x}_{v}=\mathbf{x}_{w}$, $\mathcal{N}^{1}(v)=\mathcal{N}^{2}(w)$ and $\mathcal{N}^{2}(v)=\mathcal{N}^{1}(w)$. Now, assuming $\mathcal{N}^{1}(v)\neq\mathcal{N}^{2}(v)$, the representation of $w$ can be rewritten as
$\begin{split}\mathbf{z}_{w}=\mathbf{r}_{v}\otimes\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{2}(v)}\mathbf{r}_{j}\Big{)}\otimes\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{1}(v)}\mathbf{r}_{m}\Big{)}\end{split}$
Since Binding is both associative and commutative, the latent representations $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ become equal, despite their 1-hop and 2-hop neighbors being different. This poses a problem is terms of discriminative power of the model, leading to a lack of expressive power in our HD-algorithm. To address this issue, we utilize the Rotation $\Pi$ operator as it protects/preserves information. To be more specific, with the incorporation of the rotation/permutation operator, the latent representation for nodes $v\in\mathcal{V}$ and $w\in\mathcal{V}$ changes to
@@ 3.2.3 Improving injectivity via Rotation
3.3 Node Label Prediction in HDGL
@@
With the procedure by which HDGL constructs the HD representation for each of the nodes in a graph, we turn to the task of using the resulting HD node representation to predict the node labels using $\mathbf{z}_{v}$. Here, we follow the standard approach to predicting class labels of HD encoded objects Ge and Parhi (2020). Specifically, during the learning phase, we Bundle the collection of HD encodings of nodes belonging to each class yielding a HD Class Hyper-Vector for each of the classes. During the inference phase, given the HD vectors for each of the classes, we predict the label of an unlabeled node $u$ as simply the label associated with the class hyper-vector that is, among all class HD class vectors is closest to the HD vector $\mathbf{z}_{u}$, the encoding of $u$. More precisely let $\forall\ell\in\{1,\cdots,L\}$, $\mathcal{T}_{\ell}\subset\mathcal{V}$ be the subset of nodes that make up the training set with label $\ell$ and $\mathcal{V}_{\text{test}}\subset\mathcal{V}$ be the subset of unlabeled nodes in the graph. The class HD are constructed for each class as follows:
$\mathbf{c}_{\ell}=\mathop{\oplus}_{i\in\mathcal{T}_{\ell}}\mathbf{z}_{i}\quad;\;\forall\ell\in\{1,\cdots,L\}$
resulting in a set of class HD vectors $\{\mathbf{c}_{1},\cdots,\mathbf{c}_{L}\}$. Let $u\in\mathcal{V}_{\text{test}}$ be a node for which the class label is to be predicted. We predict the class label for $u$ using its HD representation and the HD representations of each of the classes as follows:
$\hat{y}_{u}=\mathop{\operatorname{argmin}}_{\ell\in\{1,\cdots,L\}}d_{H}[\mathbf{z}_{u},\mathbf{c}_{\ell}].$

### v2

@@ 4 Hyper-Dimensional Graph Learner
The proposed HD learner carries out 3 major steps: (i). Map Node features $\mathbf{x}_{v}\in\mathbb{R}^{d}$ to a intermediate HD representation $\mathbf{r}_{v}\in\{0,1\}^{\beta}$ $\forall v\in\mathcal{V}$. (ii). Construct a latent HD representation $\mathbf{z}_{v}\in\{0,1\}^{\beta}$ using $\mathbf{r}_{v}$ and $\mathcal{N}^{k}(v)$. (iii). Use $\mathbf{z}_{v}$ to predict $y_{v}$ or predict links/edges in the graph. We proceed to describe each of these steps in detail.
Figure 1: An example mapping node features to a binary space using Random Hyperplanes. A point receives a value of 1 when on the left side of a hyperplane and 0 when on the right side. Utilizing multiple hyperplanes results in a binary sketch for each point.
@@ 4.1 Mapping Node Features to HD-space
The first crucial step is to map node features from $\mathbb{R}^{d}$ space to $\{0,1\}^{\beta}$ HD space such that $\beta>>d$. We do this using random hyperplane tessellations [12], a type of Locality Sensitive Hashing [21, 7]. The key idea is to generate $\beta$ random hyperplanes in $\mathbb{R}^{d}$ and set the corresponding co-ordinate of the HD representation of a node based on whether the its feature vector in $\mathbb{R}^{d}$ lies on the positive or the negative side of the hyperplane. With $\beta$ such hyper-planes in hand, we can generate an $n$ dimensional binary sketch for each node. More precisely, let $\mathbf{A}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{\beta}$ be a matrix with each row $\mathbf{a}_{1}^{T},\cdots\mathbf{a}_{\beta}^{T}$ drawn from $\mathcal{N}(0,I_{d})$ and $\gamma$ be uniformly distributed on $[-\lambda,\lambda]$. For conciseness, we represent $\beta$ samples of $\gamma$ as $\Gamma\sim[-\lambda,\lambda]^{\beta}$. Using $\mathbf{A}$ and $\Gamma$, a randomized binary sketch $\mathbf{r}_{v}$ for a node $v\in\mathcal{V}$ is given by $\mathbf{r}_{v}=\text{sign}(\mathbf{A}\mathbf{x}_{v}+\Gamma)$. The primary purpose of adding the $\Gamma$ offset is to ensure that the not all sampled hyperplanes pass through the origin. When dealing with evenly spread points, typically in the context of real-valued vectors, setting $\Gamma$ to zero is often appropriate.
@@ 4.2 Neighborhood Aggregation Scheme
Under the simplifying assumption that nodes that have identical multi-sets of features over their $k$-hop neighborhoods are likely to also have matching local topologies, we approximate the injectivity of the embedding $\mathbf{z}_{v}$ which maps each node to its HD representation by aggregating features of nodes over a $k$-hop neighborhood of $v$ as follows: $\mathbf{z}_{v}=\phi\Big{(}\mathbf{r}_{v},\varphi_{1}(\{\mathbf{r}_{j};j\in\mathcal{N}^{1}(v)\}),\ldots,\varphi_{K}(\{\mathbf{r}_{m};m\in\mathcal{N}^{K}(v)\})\Big{)}$ where $\mathbf{z}_{v}$ denotes the HD embedding of node $v$ and $\phi(\cdot)$ and $\varphi_{k}(\cdot);k\in\{1,\cdots,K\}$ are injective functions.
Bundling for generalizing across similar nodes. We propose that for a node $v\in\mathcal{V}$; the functions $\varphi_{k}(\cdot)$ should be the Bundling operation $\oplus$ since, due to its permutation invariance, allows us to represent unordered sets (or multi-sets). Hence, we propose to use Bundling to realize $\varphi_{k}$. More precisely, $\varphi_{k}\Big{(}\{\mathbf{r}_{j};j\in\mathcal{N}^{k}(v)\}\Big{)}=\mathop{\oplus}_{j\in\mathcal{N}^{k}(v)}\mathbf{r}_{j}$ where $\mathop{\oplus}_{j\in\mathcal{N}^{k}(v)}\mathbf{r}_{j}$ denotes binding the HD representations of all of the $k$-hop neighbors of $v$. This operation is analogous to finding a high-dimensional vector that is the most similar to all of the input HD-vectors, akin to how we find the mean in Euclidean spaces and like the mean, it is not injective. However, as noted by [59], GNNs that use the mean to aggregate features over node neighborhoods [26, 18] perform quite well in practice because the mean operation becomes increasingly injective when graphs consist of diverse node features, which is often the case in real-world graphs.
@@
Binding for linking node with its neighborhood. Having incorporated the Bundling operation in our framework by assigning it to the function $\varphi_{k}(\cdot)$, the remaining task involves identifying a suitable operation for the function $\phi(\cdot)$, for coupling the representation of a node obtained using bundling (see above) with the representations of nodes in its neighborhood.
@@
It is easy to see that if all the terms in $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ are similar, i.e., $v$ and $w$ have similar node features, and similar 1-hop and 2-hop neighborhoods, then $\mathbf{z}_{v}\approx\mathbf{z}_{w}$ as desired. On the other hand, if the encodings of the features of the 2-hop neighbors of both $v$ and $w$ differ, it results in the dissimilarity between $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ a desired outcome. When not only the 2-hop neighbors but also the 1-hop neighbors of both $v$ and $w$ are dissimilar, the dissimilarity between $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ becomes even greater. The key property that makes binding attractive for realizing the function $\phi(\cdot)$ is its reflectivity, i.e., the distance between two HD vectors remains unchanged when both of them are bound to the same third vector.
@@
Dataset Nodes Edges Features Feature Type Classes Label Rate Edge Density
CORA 2708 5429 1433 $\{0,1\}$ 7 0.052 0.0004
CiteSeer 3327 4732 3703 $\{0,1\}$ 6 0.036 0.0004
PubMed 19717 44338 500 $(0,1)$ 3 0.003 0.0001
BlogCatalog 5196 171743 8189 $\mathbb{Z}^{+}$ 6 0.023 0.0127
Coauthor CS 18333 81894 6805 $\mathbb{Z}^{+}$ 15 0.016 0.0001
Coauthor Physics 34493 247962 8415 $\mathbb{Z}^{+}$ 5 0.002 0.0001
DBLP 17725 105781 768 $\mathbb{R}$ 4 0.004 0.0006
Improving injectivity via Rotation. Because Binding is associative and commutative, it can adversly impact the injectivity and consequently, therepresentational power of our HD model. For example, suppose $v\in\mathcal{V}$ and $w\in\mathcal{V}$ are two nodes in a graph with encoding obtained using HD mapping that uses Bundling to realize $\varphi_{k}$ and binding to realize $\phi(\cdot)$ and $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ are computed as per Eq. (1). Consider the scenario in which $\mathbf{x}_{v}=\mathbf{x}_{w}$, $\mathcal{N}^{1}(v)=\mathcal{N}^{2}(w)$ and $\mathcal{N}^{2}(v)=\mathcal{N}^{1}(w)$. Now, assuming $\mathcal{N}^{1}(v)\neq\mathcal{N}^{2}(v)$, the representation of $w$ can be rewritten as $\mathbf{z}_{w}=\mathbf{r}_{v}\otimes\Big{(}\mathop{\oplus}_{j\in\mathcal{N}^{2}(v)}\mathbf{r}_{j}\Big{)}\otimes\Big{(}\mathop{\oplus}_{m\in\mathcal{N}^{1}(v)}\mathbf{r}_{m}\Big{)}$. Since Binding is both associative and commutative, the latent representations $\mathbf{z}_{v}$ and $\mathbf{z}_{w}$ become equal, despite their 1-hop and 2-hop neighbors being different. This poses a problem is terms of discriminative power of the model, leading to a lack of expressive power in our HD-algorithm. To address this issue, we utilize the Rotation $\Pi$ operator as it protects/preserves information. To be more specific, with the incorporation of the rotation/permutation operator, the latent representation for nodes $v\in\mathcal{V}$ and $w\in\mathcal{V}$ changes to
@@
4.3 Node Label Prediction using HDGL
@@
With the procedure by which HDGL constructs the HD representation for each of the nodes in a graph, we turn to the task of using the resulting HD node representation to predict the node labels using $\mathbf{z}_{v}$. Here, we follow the standard approach to predicting class labels of HD encoded objects [15]. Specifically, during the learning phase, we Bundle the collection of HD encodings of nodes belonging to each class yielding a HD Class Hyper-Vector for each of the classes. During the inference phase, given the HD vectors for each of the classes, we predict the label of an unlabeled node $u$ as simply the label associated with the class hyper-vector that is, among all class HD class vectors is closest to the HD vector $\mathbf{z}_{u}$, the encoding of $u$. More precisely let $\forall\ell\in\{1,\cdots,L\}$, $\mathcal{T}_{\ell}\subset\mathcal{V}$ be the subset of nodes that make up the training set with label $\ell$ and $\mathcal{V}_{\text{test}}\subset\mathcal{V}$ be the subset of unlabeled nodes in the graph. The class HD are constructed for each class as follows: $\mathbf{c}_{\ell}=\mathop{\oplus}_{i\in\mathcal{T}_{\ell}}\mathbf{z}_{i}\quad;\;\forall\ell\in\{1,\cdots,L\}$, resulting in a set of class HD vectors $\{\mathbf{c}_{1},\cdots,\mathbf{c}_{L}\}$. Let $u\in\mathcal{V}_{\text{test}}$ be a node for which the class label is to be predicted. We predict the class label for $u$ using its HD representation and the HD representations of each of the classes as follows: $\hat{y}_{u}=\mathop{\operatorname{argmin}}_{\ell\in\{1,\cdots,L\}}d_{H}[\mathbf{z}_{u},\mathbf{c}_{\ell}]$.
4.4 Link Prediction Using HDGL
Section 4.1 and 4.2 outline a technique for constructing node representations optimized for node classification. Here we show that the same node representation can be used to support the transductive link prediction task.
Kanerva [22, 23] describes a method for encoding entries of a dataset. Let $\mathbf{a}_{1}$, $\mathbf{a}_{2}$, $\mathbf{a}_{3}$ be some randomly sampled HD vectors representing data entry attributes (eg. weight, height and gender). Let $\mathbf{\alpha}$, $\mathbf{\beta}$ and $\mathbf{\gamma}$ be values for the corresponding attributes represented in HD format. Let $\mathbf{r}$ be a HD vector such that: $\mathbf{r}=(\mathbf{a}_{1}\otimes\mathbf{\alpha})\oplus(\mathbf{a}_{2}\otimes\mathbf{\beta})\oplus(\mathbf{a}_{3}\otimes\mathbf{\gamma}).$ From the properties of HD-computing [22], it is possible to retrieve values for a corresponding attribute using the Binding $\otimes$ operator. Specifically, $\mathbf{\alpha}\approx\mathbf{r}\otimes\mathbf{a}_{1};\quad\mathbf{\beta}\approx\mathbf{r}\otimes\mathbf{a}_{2};\quad\mathbf{\gamma}\approx\mathbf{r}\otimes\mathbf{a}_{3};$
We apply the same concept to link prediction. Let $\mathbf{e}^{+}$ be a hypervector storing information about all or some of the edges present in the graph, and let $\mathbf{e}^{-}$ be a hypervector representing information about some of the non-edges in the graph. Formally: $\mathbf{e}^{+}=\mathop{\oplus}_{(u,v)\in\mathcal{E}}(\mathbf{z}_{u}\otimes\mathbf{z}_{v});\quad\mathbf{e}^{-}=\mathop{\oplus}_{(u,v)\notin\mathcal{E}}(\mathbf{z}_{u}\otimes\mathbf{z}_{v}).$ Let $\mathbf{Z}$ have $N$ rows, storing latent representations of all the nodes in the graph. That is, a row in $\mathbf{Z}$ is represented by $\mathbf{z}_{u}$, where $u\in\mathcal{V}$. Now that we have encoded edges and non-edges in the graph, the goal is to find new or missing links in the graph. This can also be thought of as reconstructing/generating the adjacency matrix.
To find possible new links we Bind $\otimes$ the edge information hypervector $\mathbf{e}^{+}$ and the latent representation of all the nodes in the graph. This results in a new matrix $\mathbf{Z}^{+}$ where each row is $\mathbf{z}_{u}\otimes\mathbf{e}^{+}$ where $u\in\mathcal{V}$. $\mathbf{Z}^{+}$ represents the possible node representations of the target nodes in a link, with the corresponding source node representation being $\mathbf{Z}$. We follow the same procedure for the non-edge hypervector, resulting in $\mathbf{Z}^{-}$. Precisely:
$\mathbf{Z}^{+}$ $=\begin{bmatrix}\mathbf{z}_{1}\otimes\mathbf{e}^{+}\\ \vdots\\ \mathbf{z}_{N}\otimes\mathbf{e}^{+}\end{bmatrix}$ $\mathbf{Z}^{-}$ $=\begin{bmatrix}\mathbf{z}_{1}\otimes\mathbf{e}^{-}\\ \vdots\\ \mathbf{z}_{N}\otimes\mathbf{e}^{-}\end{bmatrix}$
Now, we find pairwise normalized distances between rows of $\mathbf{Z}^{+}$ and rows of $\mathbf{Z}$. This results in $\mathbf{D}^{+}=dist(\mathbf{Z}^{+},\mathbf{Z})\in[0,1]^{N\times N}$ , a matrix representing distances. If entry $(i,j)$ is close to zero, it indicates that the distance between row $i$ of $\mathbf{Z}^{+}$ is close to row $j$ in $\mathbf{Z}$, suggesting that $(i,j)$ might be a possible link. We perform the same process for $\mathbf{Z}^{-}$, resulting in the $\mathbf{D}^{-}=dist(\mathbf{Z}^{-},\mathbf{Z})\in[0,1]^{N\times N}$ matrix. However, an entry close to zero in $\mathbf{D}^{-}$ indicates a high possibility of not having an edge between the corresponding nodes.
We now construct a prediction of the adjacency matrix represented by $\hat{\mathbf{A}}$, where entries $\hat{A}_{ij}\in[0,1]$. To construct the final reconstructed or predicted adjacency matrix of probabilities (based on distances), which factors in both the distances of having, or not having an edge using $\mathbf{Z}^{+}$ and $\mathbf{Z}^{-}$, we compute:
$\hat{A}_{ij}=\begin{cases}\sigma\big{(}(1-D^{+}_{ij})+D^{-}_{ij}\big{)}&\text{if }D^{+}_{ij}<D^{-}_{ij}\\ \sigma\big{(}D^{+}_{ij}-(1-D^{-}_{ij})\big{)}&\text{else}\end{cases}$
where $\sigma(\cdot)$ is the sigmoid function. The rationale behind this operation lies in the comparison between $D^{+}_{ij}$ and $D^{-}_{ij}$. When $D^{+}_{ij}<D^{-}_{ij}$, it indicates a higher likelihood of a connection between nodes $i$ and $j$ according to the distance metric. Hence, we apply the operation $\sigma((1-D^{+}_{ij})+D^{-}_{ij})$ to obtain a value closer to one. We use the same logic when $D^{+}_{ij}\geq D^{-}_{ij}$ to obtain a value close to zero.